ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 49, 34, 93 Vasilieva, Olga, Universidad del Valle, Colombia Optimality Conditions of Singular Controls The class of problems under consideration is restricted to the optimal control problems with boundary conditions. The second-order optimality conditions of singular controls were obtained in [1] for the dynamic system with linear boundary conditions. These conditions were put into the basis of optimal design algorithms and which has resulted in so=called “improvement procedure“ allowing to continue the solution process when singular controls appear. The results of [1] are extended for the case of nonlinear boundary conditions. The application of these optimality conditions to singular optimal control design will also be presented. References: O. O. Vasilieva, and K. Mizukami, Optimality Criterion for Singular Controller: Linear Boundary Conditions, J. Math. Analysis and Applications, 213 (1997), pp.620-641.